OpenStudy (anonymous):

The line with m = 4 and intersecting 2x -4y = 8 at x = 6.

OpenStudy (callisto):

And the question is.... ?

OpenStudy (anonymous):

I need to find the line with m = 4 and intersecting 2x -4y = 8 at x = 6.

OpenStudy (callisto):

First, find the coordinates of the intersection point. Since the intersection point is on the line 2x -4y = 8 at x = 6, put x=6 into 2x -4y = 8. Solve y. 2(6) -4y = 8 12 - 4y = 8 y = .... After finding the point, you can write the equation.\[\frac{y-y_1}{x-6} = m\], where m is given in the question and \(y_1\) is the value you've found above.

OpenStudy (anonymous):

ok thanks for the steps, should be a breeze now

OpenStudy (callisto):

welcome